### v2

# Quantizing Geodesics in Kähler and Sasaki Geometry

Quantizing Geodesics in Kähler and Sasaki Geometry
Gilles Courtois, Eleonora Di Nezza, Thomas Franzinetti
Résumé
The space of Kähler potentials can be quantized through the classical Fubini–Study map, relating infinite-dimensional geometric structures to finite-dimensional symmetric spaces. We prove (exactly) when the Fubini–Study image of a geodesic line in the space of positive definite Hermitian matrices gives rise to a quasi-geodesic in the space of Kähler potentials. Furthermore, we introduce a quantization procedure for geodesics between potentials on normal Kähler varieties and show how this construction extends to the Sasaki setting.
2020 Mathematics Subject Classification : 32Q15, 53C25, 58B20, 53D50, 54E40.
1 Introduction
Let $X$ be a compact Kähler manifold and $\omega$ be a Kähler form of complex dimension $n$. The space of Kähler potentials :
$\mathcal{H}_{\omega}=\left\{u\in\mathcal{C}^{\infty}(X,\mathbb{R})\,|\,\omega_{u}:=\omega+dd^{c}u>0\right\}$ (1)
plays a central role in Kähler geometry. It is an infinite-dimensional manifold whose tangent space at each point can be identified with $C^{\infty}(X)$. At each $u\in\mathcal{H}_{\omega}$, the Mabuchi $L^{2}$-metric [28] on the tangent space at $u$ is
$\left\langle\psi_{1},\psi_{2}\right\rangle_{u}=\frac{1}{V}\int_{X}\psi_{1}\psi_{2}\,\omega_{u}^{n}\qquad\psi_{1},\psi_{2}\in T_{u}\mathcal{H}_{\omega}\simeq\mathcal{C}^{\infty}(X),$ (2)
where $V:=\int_{X}\omega^{n}>0$.
More generally, one may equip each tangent space with $L^{p}$-Finsler structures
$\|\psi\|_{u}=\left(\frac{1}{V}\int_{X}|\psi|^{p}\omega_{u}^{n}\right)^{1/p}\qquad\psi\in T_{u}\mathcal{H}_{\omega}\simeq\mathcal{C}^{\infty}(X),$ (3)
leading to a family of distances $d_{p}$ ($p\geq 1$), and a well defined notion of geodesics between two elements of $\mathcal{H}_{\omega}$. In fact, $d_{p}$ is defined as the length pseudometric associated to the Finsler $L^{p}$-norm, but one nontrivial point is that it is a genuine metric (i.e. separates points). Chen and his collaborators worked intensively in this direction [7, 6, 10, 8, 9] proving in particular that this infinite dimensional space is a path metric space (the metric is the infimum of the length of paths) with $\mathcal{C}^{1,\overline{1}}$ geodesics. Later Chu-Tosatti-Weinkove [11] has established that geodesics are $\mathcal{C}^{1,1}$-regular, which is known to be the optimal regularity thanks to examples of Darvas, Lempert, Vivas [12, 16, 27]. In particular, geodesics in general leave the space $\mathcal{H}_{\omega}$ ; we will refer to those as weak geodesics.
As shown by Chen [7] and [13] the distance $d_{p}$ between two smooth potentials $u_{0},u_{1}$ is realized by the geodesic $u_{t}$ :
$d_{p}(u_{0},u_{1})^{p}=\int_{X}|\dot{u}_{t}|^{p}(\omega+dd^{c}u_{t})^{n},\;\forall t\in[0,1].$ (4)
The understanding of the geometry and topology of $\mathcal{H}_{\omega}$ is crucial for the study of special metrics in Kähler geometry (such as Kähler-Einstein, constant scalar curvature, extremal metrics…). A powerful method to study this infinite-dimensional space is the so called Kähler quantization, which is a process to approximate the infinite dimensional space $\mathcal{H}_{\omega}$ and its geometry by finite dimensional negatively curved manifolds. Such a process has been at the core of intense research over the past decades [19, 29, 32, 9, 1, 4, 17].
The "quantization process" requires the Kähler manifold to be projective. More precisely, we consider $(X,\omega)$ a compact Kähler manifold which carries a holomorphic line bundle $L\rightarrow X$ with a Hermitian metric whose curvature is $\omega$. In this setting, $X$ comes with a family of embeddings into projective spaces :
$\Phi_{k}:X\hookrightarrow\mathbb{P}\left(H^{0}(X,L^{k})\right),$
where $H^{0}(X,L^{k})$ is the finite dimensional space of holomorphic sections of the $k$-th tensor power of $L$. By construction $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}$ lies in the same cohomology class as $\omega$}, hence $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}=\omega+dd^{c}u$, for some $u$ in $\mathcal{H}_{\omega}$. Such smooth function is in the image of the Fubini-Study map
$\operatorname{FS}_{k}:\operatorname{Herm}^{++}(H^{0}(X,L^{k}))\longrightarrow\mathcal{H}_{\omega}.$
We refer to Section 2 for a definition. Here
$\hat{\mathcal{H}}(L^{k}):=\operatorname{Herm}^{++}(H^{0}(X,L^{k}))=\operatorname{GL}(N_{k},\mathbb{C})/\operatorname{U}(N_{k},\mathbb{C})$
is the symmetric Riemannian manifold which can be interpreted as the set of Hermitian scalar products on the (finite $N_{k}$-dimensional) space of holomorphic sections $H^{0}(X,L^{k})$.
Let $\hat{d}$ be the distance on $\hat{\mathcal{H}}(L^{k})$. Our first result states the following :
Theorem 1.
The map $\operatorname{FS}_{k}:(\hat{\mathcal{H}}(L^{k}),\hat{d})\rightarrow(\mathcal{H}_{\omega},d_{p})$ is Lipschitz. More precisely,
$d_{p}(\operatorname{FS}_{k}(H_{1}),\operatorname{FS}_{k}(H_{2}))\leq C_{k}\hat{d}(H_{1},H_{2}),\quad\forall H_{1},H_{2}\in\hat{\mathcal{H}}(L^{k}).$
Here, $C_{k}$ is an explicit constant depending only on $k,X$, $\omega$ and $p$.
More generally, given a geodesic line $\mathbb{R}\ni t\mapsto H(t)$ in $\hat{\mathcal{H}}(L^{k})$, it is natural to investigate how closely the paths $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ approximate geodesics. It is known that $\operatorname{FS}_{k}(H(t))\in\mathcal{H}_{\omega}$ for any fixed $t$ and that for any geodesic segment $0,1]\ni t\mapsto H(t)\in\hat{\mathcal{H}}(L^{k})$, the path $\operatorname{FS}_{k}(H(t))$ lies below the geodesic between $\operatorname{FS}_{k}(H_{0})$ and $\operatorname{FS}_{k}(H_{1})$ [29] (i.e. $\operatorname{FS}_{k}(H(t))$ is a subgeodesic). Moreover, such subgeodesic converges in the $\mathcal{C}^{0}(X)$-topology to the geodesic segment in $\mathcal{H}_{\omega}$.
It is well known that a geodesic line in $\hat{\mathcal{H}}(L^{k})$ passing through $H_{0}=PP^{*}$, $P\in GL(N,\mathbb{C})$, reads as $H(t)=P\exp(tA)P^{*}$. We denote by $\lambda_{1}\leq\cdots\leq\lambda_{N_{k}}$ the eigenvalues of $A$. Our next result concerns the globally defined path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ :
Theorem 2.
Let $\mathbb{R}\ni t\mapsto H(t)$ be a geodesic in $\hat{\mathcal{H}}(L^{k})$, parametrized with constant speed. Assume $A$ is either definite or $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$. Then the path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ is a quasi-geodesic in $\mathcal{H}_{\omega}$.
Recall that a quasi-geodesic in a metric space $(\mathcal{M},d_{\mathcal{M}})$ is a path $\mathbb{R}\ni t\mapsto\gamma_{t}\in\mathcal{M}$ satisfying :
$\forall t,u\in\mathbb{R},\quad C^{-1}\left|t-u\right|-K\leq d_{\mathcal{M}}(\gamma_{t},\gamma_{u})\leq C\left|t-u\right|+K,$
for some $C\geq 1$ and $K\geq 0$.
We also emphasize that the assumptions on the eigenvalues of $A$ are necessary as the example in Remark 2.9 shows.
It is also worth mentioning that Phong–Sturm [30], Ross–Witt Nyström [31], and Finski [21] (in the more general big setting) show that, given a sequence (in $k$) of geodesic rays on $H^{0}(X,L^{k})$ constructed from carefully chosen norms on $H^{0}(X,L^{k})$ and appropriate speeds, the resulting sequence of quasi-geodesics converges, as $k\rightarrow+\infty$, to a weak geodesic ray. An interesting open question is whether, and under what conditions, the quasi-geodesic $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ (defined on the entire real line) converges to a weak geodesic line.
Motivated by the study of the space of Kähler potentials on singular varieties [5, 14, 18], we consider the space of potentials with respect to a big form. More precisely, as before we work on $(X,\omega)$ a compact smooth Kähler manifold and we fix another smooth $(1,1)$-form $\theta$, which we assume to be real, closed, semi-positive and big. We say that $\theta$ is big if $\theta$} can be represented by a strictly positive current, or in other words, if there exists $u\in\operatorname{PSH}(X,\theta)$ such that $\theta+dd^{c}u\geq\varepsilon\omega$, for some $\varepsilon>0$. When $\theta$ is semi-positive this simply amounts to $\int_{X}\theta^{n}>0$.
We recall that $\operatorname{PSH}(X,\theta)$ is the set of $\theta$-plurisubharmonic functions, i.e. the set of upper semicontinuous functions $\{u:X\to\mathbb{R}\cup\{-\infty\}$ which locally write as the sum of a smooth and a plurisubharmonic function and such that $\theta_{u}:=\theta+dd^{c}u\geq 0$ in the weak sense of currents. We set
$\mathcal{H}_{\theta}:=\{u\in\operatorname{PSH}(X,\theta)\cap L^{\infty}(X)\cap C^{\infty}(\operatorname{Amp}(\theta))\;\,:\;\,\theta_{u}:=\theta+dd^{c}u>0\;{\rm{in}}\;\operatorname{Amp}(\theta)\}.$
Here $\operatorname{Amp}(\theta)$ is the open set of points $x\in X$ such that there exists $\phi\in\operatorname{PSH}(X,\theta)$ which is smooth near $x$ and with analytic singularities (that is it can be locally written : $\phi=c\log\sum_{j}\left|f_{j}\right|^{2}+f$ where the $f_{j}$’s are holomorphic, $f$ is smooth and $c$ is a constant). We also mention that $\operatorname{Amp}(\theta)$ is a Zariski dense open subset of $X$ and it only depends on the cohomology class of $\{\theta\}$. Moreover, it is shown in [3, Theorem 3.17] that $\theta$ is Kähler if and only if $\operatorname{Amp}(\theta)=X$.
In [18], the authors extended Mabuchi geometry to $\mathcal{H}_{\theta}$, defining a distance and a notion of geodesics between potentials in $\mathcal{H}_{\theta}$. Here we provide a quantization result in this more general setting generalizing Berndtsson’s approximation of geodesics :
Theorem 3.
Let $u_{0},u_{1}\in\mathcal{H}_{\theta}$ and $0,1]\ni t\rightarrow u_{t}$ be the weak geodesic joining them. Then there exists a sequence of paths $(\hat{u}_{t,\ell})_{\ell\in\mathbb{N}}$ in the image of $\frac{1}{\ell}\operatorname{FS}_{k}$, where $k$ depends on $\ell$, such that for any compact $K\subset\operatorname{Amp}(\theta)$, $\hat{u}_{t,\ell}$ converges uniformly to $u_{t}$ in $K\times[0,1]$.
We point out that the recent preprint [21, Theorem 1.5 and Proposition 5.2] provides an alternative proof of the convergence stated in Theorem C in the general big (not necessarily semi-positive) setting. As explained in [21, Remark 1.6], the uniform convergence was anticipated. Our arguments yields uniform convergence directly in the semi-positive case.
Building on the above result, we develop a corresponding quantization process in the Sasaki setting. Recall that a Sasaki manifold $(M,g)$ is an odd dimensional compact Riemannian manifold whose metric cone $C(M)=(\mathbb{R}^{*}_{+}\times M,dr^{2}+r^{2}g)$ is Kähler. In particular, $M$, which can be identified with the link $\left\{r=1\right\}\subset C(M)$, is a contact manifold with contact form $\eta=2d^{c}\log(r)$. It defines a contact bundle $\ker\eta$ on which $\frac{1}{2}d\eta$ is a transverse Kähler form. Moreover, any Sasaki manifold is endowed with a special vector field : the Reeb vector field $\xi$ which is the restriction of $J(r\partial_{r})$ to the link $\left\{r=1\right\}$. Here $J$ denotes the complex structure on the Kähler metric cone $C(M)$. The restriction $\Phi$ of $J$ to the transverse distribution $\ker\eta$ is called a transverse complex structure. Using $\xi,\eta$ and $\Phi$, one can define a Riemannian metric : $g=\eta\otimes\eta+\frac{1}{2}d\eta(\cdot,\Phi\cdot)$. We call $\mathcal{S}:=(\xi,\eta,\Phi,g)$ a Sasaki structure.
By analogy with Kähler potentials, we consider the space of potentials on a Sasaki manifold $(M,\mathcal{S})$ :
$\mathcal{H}(M,\mathcal{S})=\left\{\phi\in\mathcal{C}^{\infty}_{B}(M),\;d\eta+dd^{c}\phi>0\right\},$
where $\mathcal{C}^{\infty}_{B}(M)$ is the space of smooth basic functions (ie smooth functions which are invariant under the Reeb flow). Analogously to the Kähler setting, the $d^{c}$-operator, acting on basic functions, is defined as $\frac{i}{2}(\overline{\partial}_{B}-\partial_{B})$ where the $\partial_{B}$ and $\overline{\partial}_{B}$ operator are defined in [boyer_sasakian_2007]. We omit the subscript $B$ for simplicity. This infinite dimensional space, whose tangent space at any $u\in\mathcal{H}(M,\mathcal{S})$ is identified with $\mathcal{C}_{B}^{\infty}(M)$, is endowed with a Riemannian structure, analogue of the Mabuchi metric [23, 24] :
$\left\langle\psi_{1},\psi_{2}\right\rangle_{u}=\frac{1}{\int_{M}\eta\wedge d\eta^{n}}\int_{M}(\psi_{1}\psi_{2})\;\eta\wedge(d\eta+dd^{c}u)^{n}.$
Here and in the sequel, $2n+1$ is the real dimension of $M$.
Guan and Zhang [23, 24] proved the existence of $\mathcal{C}^{1,\overline{1}}$ geodesics using a Monge-Ampère type re-formulation for the geodesic equation. They also showed that the Riemannian structure on the tangent space of $\mathcal{H}(M,\mathcal{S})$ induces a distance on $\mathcal{H}(M,\mathcal{S})$.
We will make use of the results in the Kähler setting to study quantization for regular and quasi-regular Sasaki manifolds. Indeed, recall that if a Sasaki manifold is quasi-regular, then the space of orbits of the Reeb vector field is a Kähler orbifold. If it is regular, then the space of orbits is a Kähler manifold.
If $(M,\mathcal{S})$ is a regular or quasi-regular Sasaki manifold, we let $\pi:M\rightarrow X$ be the projection onto the space of orbits under the Reeb vector field. We let $M^{\text{reg}}:=\pi^{-1}(X^{\text{reg}})$ be the set of points in $M$ which lie in a regular orbit under the Reeb vector field (if $M$ is regular, then $M^{\text{reg}}=M$). In addition, we consider $\hat{\mathcal{H}}(L^{k})$, the space of Hermitian metrics on the finite dimensional space of holomorphic sections of a certain power of a line bundle $L$ over a resolution of singularities of $X$ (see Section 4). In this setting we have :
Theorem 4.
Let $u_{0},u_{1}\in\mathcal{H}(M,\mathcal{S})$ and let $u_{t}$ be the weak geodesic joining $u_{0}$ and $u_{1}$. There exists a sequence of paths $(\hat{u}_{t,\ell})_{\ell\in\mathbb{N}}$, which are images of geodesics in $\hat{\mathcal{H}}(L^{k})$, such that for any compact set $K\subset M^{\text{reg}}$, $\hat{u}_{t,\ell}$ converges uniformly to $u_{t}$ in $K\times[0,1]$ .
Organization of the paper.
Section 2 recalls the geometry of Hermitian symmetric spaces and the definition of the Fubini–Study map, and proves Theorems 1 and 2. Section 3 establishes the quantization result in the semi-positive and big setting (Theorem 3 ). Section 4 develops the Sasaki framework and proves Theorem 4.
Acknowledgements
The second author is supported by the ERC grant SiGMA (No. 101125012, PI : Eleonora Di Nezza). She also gratefully acknowledges the University of Melbourne (MATRIX campus) for providing a welcoming and stimulating environment during the research program “Analytic and Geometric Methods on Complex Manifolds" in which this project was completed. Moreover, we thank Sébastien Boucksom, Tamas Darvas, Siarhei Finski and Mattias Jonsson for helpful discussions and comments on a previous draft of the note.
2 The Fubini-Study map $\operatorname{FS}_{k}$
After a brief review over the geometry of the set of Hermitian matrices, we will recall the definitions of the main two maps $\operatorname{FS}_{k}$ and $\operatorname{Hilb}_{k}$ which are at the core of Kähler quantization.
The space of Hermitian Matrices
We furnish the space of Hermitian matrices with a structure of symmetric Riemannian manifold. We are particularly interested in the explicit description of geodesics.
Let $\operatorname{Herm}^{++}(N,\mathbb{C})$ be the space of definite positive Hermitian matrices of size $N\times N$. We furnish $\operatorname{Herm}^{++}(N,\mathbb{C})$ with a Riemannian metric :
$\left\langle H_{1},H_{2}\right\rangle_{H}:=\frac{1}{N}\operatorname{Tr}(H^{-1}H_{1}H^{-1}H_{2}),$ (5)
where $H\in\operatorname{Herm}^{++}(N,\mathbb{C})$ and $H_{1},H_{2}\in T_{H}{\operatorname{Herm}^{++}(N,\mathbb{C})}\simeq\operatorname{Herm}(N,\mathbb{C})$. The scaling factor $\frac{1}{N}$ will appear to be nice in our setting since the dimension of the matrices that we will consider will be bigger and bigger.
Let us recall that $\text{Herm}^{++}(N,\mathbb{C})$ is a symmetric space of non positive curvature, [25]. For $P\in GL(N,\mathbb{C})$, we denote $P^{*}$ the adjoint of $P$. The action of $GL(N,\mathbb{C})$ on $\text{Herm}^{++}(N,\mathbb{C})$ defined for $P\in GL(N,\mathbb{C})$ and $H\in\text{Herm}^{++}(N,\mathbb{C})$ by $P.H:=PHP^{*}$ is a transitive isometric action. In particular every $H\in\text{Herm}^{++}(N,\mathbb{C})$ can be written as $P.Id=PP^{*}$ for some $P\in GL(N,\mathbb{C})$. Hence, $\text{Herm}^{++}(N,\mathbb{C})$ may be identified with $GL(N,\mathbb{C})/\penalty 50U(N,\mathbb{C})$, where $U(N,\mathbb{C})$ is the stabilizer of $Id$. A path $t\mapsto H(t)\in\operatorname{Herm}^{++}(N,\mathbb{C})$ is a geodesic with respect to the metric (5) if and only if $\ddot{H}=\dot{H}H^{-1}\dot{H}$, or equivalently : $\frac{d}{dt}\left(H^{-1}\dot{H}\right)=0.$
As a consequence, geodesics $H(t)$ in $\operatorname{Herm}^{++}(N,\mathbb{C})$ such that $H(0)=H_{0}$ all are of the form :
$\mathbb{R}\ni t\mapsto H(t)=P\exp(tA)P^{*},$ (6)
where $H_{0}=PP^{*}$ with $P\in GL(N,\mathbb{C})$ and $PAP^{*}\in\operatorname{Herm}(N,\mathbb{C})$ is the initial direction. Notice that the above geodesics have constant speed equal to $\sqrt{\frac{1}{N}\text{Tr}(A^{2})}$, in particular, for any $s<t\in\mathbb{R}$, the riemannian distance between $H(s)$ and $H(t)$ is equal to
$\hat{d}(H(s),H(t))=\sqrt{\frac{1}{N}\operatorname{Tr}\left(A^{2}\right)}\left|t-s\right|.$ (7)
The Kähler Quantization
Consider $(L,h_{0})\rightarrow X$ a positive line bundle whose curvature is $\omega$. Observe that, given any $u\in\mathcal{H}_{\omega}$, $h_{0}e^{-u}$ is an other positive hermitian metric on $L$. By abuse of notation we will still denote by $h_{0}$ the induced hermitian metric on the space of holomorphic sections $H^{0}(X,L^{k})$, defined as
$h_{0}(s,s^{\prime}):=\int_{X}h_{0}(s(x),s^{\prime}(x))\omega^{n},\qquad s,s^{\prime}\in H^{0}(X,L^{k}).$
The space $H^{0}(X,L^{k})$ is a finite dimensional space, we write $N_{k}$ its dimension. Let $\hat{\mathcal{H}}(L^{k})$ be the space of Hermitian metrics on $H^{0}(X,L^{k})$. The space $\hat{\mathcal{H}}(L^{k})$ can be identified with $\operatorname{Herm}^{++}(N_{k},\mathbb{C})$. Indeed, given $\{s_{i}\}$ an orthonormal basis for $h_{0}$, any $h\in\hat{\mathcal{H}}(L^{k})$ is associated with a unique $H\in\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ so that $H_{ij}=h(s_{i},s_{j})$. In the following, using the above identification, we will talk about the $H$-orthonormal basis of sections, instead of $h$-orthonormal.
The goal here is to use the finite dimensional spaces $\hat{\mathcal{H}}(L^{k})$ to approximate the infinite dimensional space $\mathcal{H}_{\omega}$. Such a quantization process involves two maps :
$\operatorname{Hilb}_{k}:\mathcal{H}_{\omega}\rightarrow\hat{\mathcal{H}}(L^{k}),\qquad\operatorname{FS}_{k}:\hat{\mathcal{H}}(L^{k})\rightarrow\mathcal{H}_{\omega}.$
Let us start with the definition of the map $\operatorname{Hilb}_{k}:\mathcal{H}_{\omega}\rightarrow\hat{\mathcal{H}}(L^{k})$ associating to each $u\in\mathcal{H}_{\omega}$ the Hermitian metric :
$\forall s\in H^{0}(X,L^{k}),\quad\operatorname{Hilb}_{k}(u)(s,s):=\int_{X}\left|s\right|_{(h_{0}e^{-u})^{k}}^{2}\omega^{n}=\int_{X}\left|s\right|_{h_{0}^{k}}^{2}e^{-ku}\omega^{n}.$
Remark 2.1.
Note that the volume form we use here is not the same as in the classical definition [20, p347] of the $\operatorname{Hilb}_{k}$ map.
In the opposite direction, given $H$ a Hermitian metric on $H^{0}(X,L^{k})$, we consider $\left\{s_{j}\right\}$ a $H$-orthonormal basis and we set :
$\operatorname{FS}_{k}(H)=\frac{1}{k}\log\left(\sum_{j}\left|s_{j}\right|_{h_{0}^{k}}^{2}\right)\in\mathcal{H}_{\omega}.$ (8)
We recall a fundamental result due to Berndtsson [1, Theorem 1.2] :
Theorem 2.2.
Let $u_{0},u_{1}\in\mathcal{H}_{\omega}$, and $0,1]\ni t\mapsto u_{t}$ be the weak geodesic joining $u_{0}$ and $u_{1}$. We write $H_{i}=\operatorname{Hilb}_{k}(u_{i})$ for $i=0,1$ and we let $H(t)$ be the geodesic in $\hat{\mathcal{H}}(L^{k})$ from $H_{0}$ to $H_{1}$. Then there is a constant $C>0$ such that :
$\sup_{(x,t)\in X\times[0,1]}\left|\operatorname{FS}_{k}(H(t))(x)-u_{t}(x)\right|\leq C\frac{\log k}{k}.$
In particular, the weak geodesic $u_{t}$ can be approximated in the $\mathcal{C}^{0}(X)$-topology by the quantized path $\operatorname{FS}_{k}(H(t))$.
As mentioned in Remark 2.1, our definition of the map $\operatorname{Hilb}_{k}$ is not the classical one. It follows from [1] that we can choose any volume form in the definition of $\operatorname{Hilb}_{k}$ so that the above classical approximation theorem is still valid. We refer to [17, Theorem 1.2] for more details.
Quantized geodesic lines are quasi-geodesics
In view of Theorem 2.2, the image by $\operatorname{FS}_{k}$ of the geodesic segment between $\operatorname{Hilb}_{k}(u_{0})$ and $\operatorname{Hilb}_{k}(u_{1})$ in $\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ is close, in $\mathcal{C}^{0}$-norm, to the geodesic in $\mathcal{H}_{\omega}$ between the two potentials $u_{0}$ and $u_{1}$, at least for $k$ large enough. Notice that this holds for bounded geodesic segment. We now study the image by the map $FS_{k}$ of complete geodesic lines $\mathbb{R}\rightarrow\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ in (6), keeping in mind the same question as above : are they close to geodesics lines in $\mathcal{H}_{\omega}$ ? We don’t answer this question but we prove that $FS_{k}$ maps geodesic lines in $\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ onto quasi-geodesics in $\mathcal{H}_{\omega}$. Let us recall the notion of quasi-geodesics :
Definition 2.3.
A quasi-geodesic in a metric space $(\mathcal{M},d_{\mathcal{M}})$ is a path $I\ni t\mapsto\gamma_{t}\in\mathcal{M}$ such that there exist constants $C,K>0$ such that :
$\forall t,s\in I,\quad C^{-1}\left|t-s\right|-K\leq d_{\mathcal{M}}(\gamma_{t},\gamma_{s})\leq C\left|t-s\right|+K.$
Here, $I\subset\mathbb{R}$ is an interval.
In what follows we will need the following two lemmas. The first one provides a comparison between the Mabuchi distance and some explicit analytic expression.
Lemma 2.4 ([15, Theorem 3.32]).
There exists $C=C(p)>0$ such that for any $u_{0},u_{1}\in\mathcal{H}_{\omega}$ we have :
$\frac{1}{C}d_{p}(u_{0},u_{1})^{p}\leq\frac{1}{V}\int_{X}\left|u_{0}-u_{1}\right|^{p}\left(\omega_{u_{0}}^{n}+\omega_{u_{1}}^{n}\right)\leq Cd_{p}(u_{0},u_{1})^{p}.$
The second lemma describes the variation of the potential $FS_{k}(H(t))$ along a geodesic $H(t)$ of $\hat{\mathcal{H}}(L^{k})$ identified with $\operatorname{Herm}^{++}(N_{k},\mathbb{C})$.
Lemma 2.5.
Let $A\in\operatorname{Herm}(N_{k},\mathbb{C})$ and $H_{0}=PP^{*}\in\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ and $H(t):=P\exp(tA)P^{*}$ a geodesic passing through $H(0)=H_{0}$. Then, for every $s\leq t$,
$\frac{\max(\lambda_{1},-\lambda_{N_{k}})}{k}\left|t-s\right|\leq\left|\operatorname{FS}_{k}(H(t))-\operatorname{FS}_{k}(H(s))\right|\leq\frac{\max(-\lambda_{1},\lambda_{N_{k}})}{k}\left|t-s\right|,$ (9)
where $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N_{k}}$ are the eigenvalues of $A$.
Démonstration.
Let $\{s_{j}\},j=1,...,N_{k}$, be a $H_{0}$-orthonormal basis of $\hat{\mathcal{H}}(L^{k})$. We claim that $\{s_{j}e^{-\frac{\lambda_{j}t}{2}}\}$ is an orthonormal basis for $H(t)=P\exp(tA)P^{*}$. To see this, we first observe that $\{E_{j}\}$ is an orthonormal basis for $H_{0}=PP^{*}$ w.r.t. $h_{0}:=<\cdot,\cdot>_{0}$ (i.e. $PP^{\star}E_{j},E_{k}>_{0}=\delta_{jk}$) if and only if $\{e_{j}\}:=\{P^{*}E_{j}\}$ is an orthonormal basis for $\cdot,\cdot>_{0}$.
Now, let $\{e_{j}\}$ be an orthonormal basis for $\cdot,\cdot>_{0}$ which diagonalize $A$, i.e. $e^{tA}e_{j}=e^{t\lambda_{j}}e_{j}$ then $\{e^{-{t\lambda_{j}}/{2}}E_{j}\}$ (where $E_{j}$ is defined so that $P^{*}E_{j}=e_{j}$) is an orthonormal basis for $Pe^{tA}P^{*}$. The above arguments applied to $E_{j}=s_{j}$ prove the claim.
Hence we have
$\operatorname{FS}_{k}(H(t))=\frac{1}{k}\log\left(\sum_{j}\left|s_{j}\right|_{h_{0}^{k}}^{2}e^{-t\lambda_{j}}\right).$ (10)
Let us denote
$u(t):=\operatorname{FS}_{k}(H(t)).$ (11)
We then have
$u^{\prime}(t)=-\frac{1}{k}\frac{\sum_{j}\lambda_{j}\left|s_{j}(x)\right|_{h_{0}^{k}}^{2}e^{-t\lambda_{j}}}{\sum_{j}\left|s_{j}(x)\right|_{h_{0}^{k}}^{2}e^{-t\lambda_{j}}}.$
On the other hand, for any $1\leq j\leq N_{k}$, we have $\lambda_{1}\leq\lambda_{j}\leq\lambda_{N_{k}}$, hence
$\lambda_{N_{k}}e^{-t\lambda_{j}}\left|s_{j}(x)\right|_{h_{0}^{k}}^{2}\leq-\lambda_{j}e^{-t\lambda_{j}}\left|s_{j}(x)\right|_{h_{0}^{k}}^{2}\leq-\lambda_{1}e^{-t\lambda_{j}}\left|s_{j}(x)\right|_{h_{0}^{k}}^{2}.$
Summing up the previous inequalities on $1\leq j\leq N_{k}$, we get
$\frac{\lambda_{N_{k}}}{k}\leq u^{\prime}(t)\leq-\frac{\lambda_{1}}{k},$
therefore, for any $s<t$, we have
$\frac{\lambda_{N_{k}}}{k}\leq\frac{u(t)-u(s)}{t-s}\leq-\frac{\lambda_{1}}{k}.$
We deduce
$\frac{\max(\lambda_{1},-\lambda_{N_{k}})}{k}\left|t-s\right|\leq\left|u(t)-u(s)\right|\leq\frac{\max(-\lambda_{1},\lambda_{N_{k}})}{k}\left|t-s\right|,$ (12)
which concludes the proof. ∎
Theorem 2.6.
The map $\operatorname{FS}_{k}:(\hat{\mathcal{H}}(L^{k}),\hat{d})\rightarrow(\mathcal{H},d_{p})$ is $C(k)$-Lipschitz with Lipschitz constant $C(k):=\frac{C_{p}}{k}\sqrt{N_{k}}$, where $C_{p}$ is a positive constant which depends on $p$.
Remark 2.7.
We remark that, since $N_{k}\sim k^{n}$ as $k\rightarrow+\infty$, the Lipschitz constant $C(k)$ can be chosen independently on $k$ when $n=2$. We also note that in the case of Riemann surfaces $FS_{k}$ is a contraction for $k>>1$.
We mention that Lempert [26, Theorem 1.1] proved that the map $\operatorname{FS}_{k}$ is injective.
Démonstration.
Let $H_{0},H_{1}\in\text{Herm}^{++}(N_{k},\mathbb{C})$. By (6), the geodesic segment $H(t),t\in[0,1]$ joining $H_{0}$ and $H_{1}$ writes $H(t)=P\exp(tA)P^{*}$ where $H_{0}=PP^{*}$ and $A=\log\left(P^{-1}H_{1}(P^{*})^{-1}\right)$. By (7), the distance between $H_{0}$ and $H_{1}$ is equal to $\hat{d}(H_{0},H_{1})=\sqrt{\frac{1}{N_{k}}\text{Tr}(A^{2})}$.
Denote by $u_{i}:=\operatorname{FS}_{k}(H_{i})$. From Lemma 2.4 and the right hand side of (9) we get
$d_{p}\left(\operatorname{FS}_{k}(H_{0}),\operatorname{FS}_{k}(H_{1})\right)$ $\leq\left(\frac{C_{p}}{V}\right)^{1/p}\left(\int_{X}\left|u_{0}-u_{1}\right|^{p}\left(\omega_{u_{0}}^{n}+\omega_{u_{1}}^{n}\right)\right)^{1/p}\leq C_{p}^{\prime}\,\frac{\max(\lambda_{N_{k}},-\lambda_{1})}{k}.$
Observing that $\max(\lambda_{N_{k}},-\lambda_{1})=\max(\left|\lambda_{1}\right|,\left|\lambda_{N_{k}}\right|)\leq\sqrt{\operatorname{Tr}(A^{2})}=\sqrt{N_{k}}\hat{d}(H_{0},H_{1}),$ we deduce from the previous inequality that
$d_{p}(\operatorname{FS}_{k}(H_{0}),\operatorname{FS}_{k}(H_{1}))\leq C_{p}^{\prime}\frac{\sqrt{N_{k}}}{k}\hat{d}(H_{0},H_{1}),$
which concludes the proof of Theorem 2.6. ∎
Let us consider $A\in\operatorname{Herm}(N_{k},\mathbb{C})$ such that $A\neq 0$ and $H(t):=P\exp(tA)P^{*}$ a geodesic ray in $\operatorname{Herm}^{++}(N_{k},\mathbb{C})$ passing through $H_{0}=PP^{*}$. Observe that, since $A\neq 0$, then $H(t)$ is not constant in time. The following result shows when $t\to\operatorname{FS}_{k}(H(t))$ is a quasi-geodesic in $\mathcal{H}$.
Theorem 2.8.
Let $0\neq A\in\operatorname{Herm}(N_{k},\mathbb{C})$ and $H_{0}=PP^{*}\in\hat{\mathcal{H}}(L^{k})$. Assume $A$ is either definite or $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$. Then, the image by $\operatorname{FS}_{k}$ of the geodesic $\mathbb{R}\ni t\mapsto H(t):=P\exp(tA)P^{*}$ is a quasi-geodesic in $\mathcal{H}_{\omega}$. More precisely, there exist positive constants $C_{k},D_{k}>0$ such that
$C_{k}\hat{d}(H(t),H(s))-D_{k}\leq d_{p}(\operatorname{FS}_{k}(H(t)),\operatorname{FS}_{k}(H(s)))\leq C_{p}\frac{\sqrt{N_{k}}}{k}\hat{d}(H(t),H(s)).$
Remark 2.9.
The assumption on the eigenvalues of $A$ is optimal as the following example shows. First, we observe that up to work with a large power of $L$ we can suppose that $L$ is very ample and that $H^{0}(X,L)$ has a basis $(s_{1},...,s_{N})$ so that $s_{1},...,s_{p}$ have no common zeros for some $p<N$. Now, given $H(t)=diag(1,...,1,e^{t},...,e^{t})$, ie. $\lambda_{1}=\cdots:\lambda_{p}=0<1=\lambda_{p+1}=\cdots=\lambda_{N}$, we can infer that
$\operatorname{FS}_{1}(H(t))=\log\left(\sum_{i\leq p}|s_{i}|^{2}+e^{-t}\sum_{i>p}|s_{j}|^{2}\right)$
decreases at $t$ goes to $\infty$ to $\phi=\log\sum_{i\leq p}|s_{i}|^{2}$. In particular $d_{p}(\operatorname{FS}_{1}(H(t)),\operatorname{FS}_{1}(H(0)))$ is bounded in $t$. On the other side, $\hat{d}(H(t),H(0))=\frac{1}{\sqrt{N}}t$. This implies that we can not find uniform constants (in $t$) $A,B>0$ so that $\hat{d}(H(t),H(0))\leq Ad_{p}(\operatorname{FS}_{1}(H(t)),\operatorname{FS}_{1}(H(0)))+B$.
We thank Sebastien Boucksom and Mattias Jonsson for sharing this example.
Démonstration.
By Theorem 2.6, the map $FS_{k}$ is Lipschitz, thus it remains to show the left hand side inequality of the statement. Let us denote by $\lambda_{1}\leq\cdots\leq\lambda_{N_{k}}$ the eigenvalues of $A$. Observe that if $\lambda_{1}>0$ or $\lambda_{N_{k}}<0$, then the conclusion follows straightforwardly by Lemma 2.5.
In what follows we consider the case $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$. Let $\left\{s_{j}\right\}$ of $H^{0}(X,L^{k})$ be a $H_{0}$-orthonormal basis. As we already saw in the proof of Lemma 2.5 we have
$u(t):=\operatorname{FS}_{k}(H(t))=\frac{1}{k}\log\left(\sum_{j}\left|s_{j}\right|_{h_{0}^{k}}^{2}e^{-t\lambda_{j}}\right).$
To simplify notation, in the following we simply write $\left|s_{j}\right|$ instead of $\left|s_{j}\right|_{h_{0}^{k}}^{2}$. Now, for $\tau\in\mathbb{R}^{+}$
$u^{\prime}(\tau)$ $=$ $\frac{1}{k}\frac{\sum_{\lambda_{j}<0}|\lambda_{j}|\left|s_{j}\right|^{2}e^{-\tau\lambda_{j}}-\sum_{\lambda_{i}>0}|\lambda_{i}|\left|s_{i}\right|^{2}e^{-\tau\lambda_{i}}}{\sum_{\lambda_{\ell}<0}\left|s_{\ell}\right|^{2}e^{-\tau\lambda_{\ell}}+\sum_{\lambda_{\ell}\geq 0}\left|s_{\ell}\right|^{2}e^{-\tau\lambda_{\ell}}}:=\frac{1}{k}\frac{A(\tau)-B(\tau)}{C(\tau)+D(\tau)}.$
Observe that $D(\tau)$ is bounded above independently of $\tau$ since $D(\tau)\leq\sum_{\ell}\left|s_{\ell}\right|^{2}\leq C_{1}$. Also,
$A(\tau)-B(\tau)$ $=$ $\sum_{\lambda_{j}<0}e^{-\tau\lambda_{j}}\left(|\lambda_{j}||s_{j}|^{2}-\sum_{\lambda_{i}>0}|\lambda_{i}||s_{i}|^{2}e^{-\tau(\lambda_{i}+|\lambda_{j}|)}\right)$
$\geq$ $\sum_{\lambda_{j}<0}e^{-\tau\lambda_{j}}\left(|\lambda_{j}||s_{j}|^{2}-h(\tau)\right),$
where $h:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ smooth function going to zero as $\tau\rightarrow+\infty$. Combining all the above we get that for any $t\in\mathbb{R}$
$\int_{X}u^{\prime}(\tau)\omega_{u(t)}^{n}$ $\geq$ $\frac{1}{k}\frac{\sum_{\lambda_{j}<0}e^{-\tau\lambda_{j}}\left(\int_{X}|\lambda_{j}||s_{j}|^{2}\omega_{u(t)}^{n}-h(\tau)\right)}{C_{1}\left(\sum_{\lambda_{\ell}<0}e^{-\tau\lambda_{\ell}}+1\right)}.$
It follows that
$\liminf_{\tau\rightarrow+\infty}\int_{X}u^{\prime}(\tau)\omega_{u(t)}^{n}\geq\frac{1}{C_{1}k}\inf_{\lambda_{j}<0}\int_{X}|\lambda_{j}||s_{j}|^{2}\omega_{u(t)}^{n},$
where the infimum is strictly positive since $\{s_{j}\}$ is an orthonormal basis, in particular $s_{j}$ can not be identically zero. By definition of liminf, there exists $\tau_{0}$ and $\alpha>0$ such that for any $s\in\mathbb{R}$, $\int_{X}u^{\prime}(\tau)\,\omega_{u(t)}^{n}\geq\alpha>0$ for $\tau\geq\tau_{0}$. The same arguments will give that $\int_{X}u^{\prime}(\tau)\,\omega_{u(s)}^{n}\geq\beta>0$ for $\tau\geq\tau_{0}$. Up to relabelling, we can assume that $\alpha=\min(\alpha,\beta)$.
For $\tau\in\mathbb{R}^{-}$, the same arguments, switching the role of positive and negative eigenvalues, will ensure that there exists $\tau^{\prime}_{0},\alpha^{\prime}>0$ such that $\int_{X}|u^{\prime}(\tau)|\,\omega_{u(t)}^{n}+\omega_{u(s)}^{n}\geq\alpha^{\prime}>0$ for $\tau\leq-\tau^{\prime}_{0}$. Up to relabelling again, we can assume that $\tau_{0}=\max(\tau_{0},\tau_{0}^{\prime})$ and $\alpha=\min(\alpha,\alpha^{\prime})$. We stress that in what follows we need $\alpha>0$ which means that we need to restrict to the case $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$.
Thanks to the fundamental theorem of calculus we know that for any $s<t\in\mathbb{R}$, $u(t)-u(s)=\int_{s}^{t}u^{\prime}(\tau)d\tau$. Then, using Lemma 2.4 and Jensen inequality we have
$d_{p}(u(t),u(s))^{p}$ $\geq$ $C_{p}^{-1}\int_{X}\left|\int_{s}^{t}u^{\prime}(\tau)d\tau\right|^{p}\omega_{u(t)}^{n}+\omega_{u(s)}^{n}$
$\geq$ $C_{p}^{-1}\left|\int_{s}^{t}\left(\int_{X}u^{\prime}(\tau)\,\omega_{u(t)}^{n}+\omega_{u(s)}^{n}\right)d\tau\right|^{p}$
We now distinguish five cases :
  • (i)
    $s,t>\tau_{0}$,
  • (ii)
    $s,t\in[-\tau_{0},\tau_{0}]$,
  • (iii)
    $s\in[-\tau_{0},\tau_{0}],t>\tau_{0}$,
  • (iv)
    $s<-\tau_{0},t\in[-\tau_{0},\tau_{0}]$,
  • (v)
    $s,t<-\tau_{0}$.
When $s,t>\tau_{0}$
$d_{p}(u(t),u(s))^{p}\geq C_{p}^{-1}(2\alpha)^{p}|t-s|^{p},$
hence
$d_{p}(u(t),u(s))\geq\frac{\alpha}{C_{p}^{\prime}}(t-s)=\frac{\alpha}{{C_{p}^{\prime}}}\sqrt{\frac{N_{k}}{\sum_{i}\lambda_{i}^{2}}}\,\hat{d}(H(t),H(s)).$
When $s,t\in[-\tau_{0},\tau_{0}]$ then $t-s\leq 2\tau_{0}$, hence
$d_{p}(u(t),u(s))\geq 0\geq(t-s)-2\tau_{0}=\sqrt{\frac{N_{k}}{\sum_{i}\lambda_{i}^{2}}}\,\hat{d}(H(t),H(s))-2\tau_{0},$
We now discuss the case (iii). We use the linearity of the integral and observe that
$\int_{s}^{\tau_{0}}\left(\int_{X}u^{\prime}(\tau)\,\omega_{u(t)}^{n}+\omega_{u(s)}^{n}\right)d\tau+\int_{\tau_{0}}^{t}\left(\int_{X}u^{\prime}(\tau)\,\omega_{u(t)}^{n}+\omega_{u(s)}^{n}\right)d\tau$
$\geq$ $-2C_{0}(\tau_{0}-s)+2\alpha(t-\tau_{0})=2\alpha(t-s)-(2\alpha+2C_{0})(\tau_{0}-s)$
$\geq$ $2\alpha(t-s)-(2\alpha+2C_{0})\tau_{0},$
where the first inequality follows from the fact that $|u^{\prime}(\tau)|\leq C_{0}$ on the compact interval $\tau_{0},\tau_{0}]$, since $u^{\prime}$ is a continuous function. We consider two sub-cases. Let $a_{p}=2^{p+1}(2^{p}-1)$. When $\alpha(t-s)\leq a_{p}\tau_{0}(\alpha+C_{0})$ we can conclude easily since
$d_{p}(u(t),u(s))\geq 0\geq\alpha(t-s)-a_{p}\tau_{0}(\alpha+C_{0})=\alpha\sqrt{\frac{N_{k}}{\sum_{i}\lambda_{i}^{2}}}\,\hat{d}(H(t),H(s))-a_{p}\tau_{0}(\alpha+C_{0}).$
When $\alpha(t-s)>a_{p}\tau_{0}(\alpha+C_{0})$, we observe that for each $j=0,\cdots,p-1$ we have
$\frac{c_{j,p}}{2}[\alpha(t-s)]^{p-j}>\binom{p}{j}[(\alpha+C_{0})\tau_{0}]^{b-j},$
where $c_{j,p}:=\frac{2^{j}}{2^{p}-1}$. Hence
$d_{p}(u(t),u(s))^{p}$ $\geq$ $C_{p}^{-1}(\alpha(t-s)-(\alpha+C_{0})\tau_{0})^{p}$
$\geq$ $C_{p}^{-1}\left(\sum_{j=0}^{p-1}c_{j,p}\alpha^{p}(t-s)^{p}+\sum_{j=0}^{p-1}\binom{p}{j}(-1)^{j}\alpha^{j}(t-s)^{j}(\tau_{0}(\alpha+C_{0}))^{j}\right)$
$\geq$ ${C^{\prime}}_{p}^{-1}\alpha^{p}(t-s)^{p}.$
Thus
$d_{p}(u(t),u(s))\geq{C^{\prime\prime}}_{p}^{-1}\alpha\sqrt{\frac{N_{k}}{\sum_{i}\lambda_{i}^{2}}}\,\hat{d}(H(t),H(s)).$
We can finally deduce that there exist positive constants $C_{k}=C_{k}(A,N_{k},n,\alpha,p),D_{k}=D_{k}(\alpha,p,\tau_{0})$ such that
$d_{p}(u(t),u(s))\geq C_{k}\hat{d}(H(t),H(s))-D_{k}.$
The cases (iv) and (v) follow exactly as the cases in (iii) and (i), respectively. The proof is complete. ∎
3 Quantization in the semi-positive case
Let $(X,\omega)$ be a compact Kähler manifold of complex dimension $n$ and $\theta$ be a real and closed $(1,1)$-form which is semi-positive and big. In this section we describe an approach to give a quantization result for the geodesics in the space
$\mathcal{H}_{\theta}:=\{u\in\operatorname{PSH}(X,\theta)\cap L^{\infty}(X)\cap C^{\infty}(\operatorname{Amp}(\theta))\;:\;\theta_{u}=\theta+dd^{c}u>0\;{\rm{in}}\;\operatorname{Amp}(\theta)\}.$
It was proved in [18] that $\mathcal{H}_{\theta}$ is a metric space. For $\varepsilon>0$ let $\omega_{\varepsilon}:=\theta+\varepsilon\omega>0$, and $d_{p,\omega_{\varepsilon}}$ denote the metric on $\overline{\mathcal{H}_{\omega_{\varepsilon}}}\supset\mathcal{H}_{\theta}$. In [18], the authors proved that for any $u_{0},u_{1}\in\mathcal{H}_{\theta}$, $\lim_{\varepsilon\rightarrow}d_{p,\omega_{\varepsilon}}(u_{0},u_{1})$ exists and defines a metric, denoted by $d_{p}$ as well. Moreover, for $t=0,1$, one has an explicit expression of the distance :
$d_{p}(u_{0},u_{1})=\left(\frac{1}{V}\int_{X}|\dot{u}_{t}|^{p}\theta^{n}_{u_{t}}\right)^{1/p},$
where $t\rightarrow u_{t}$ is the geodesic joining $u_{0}$ and $u_{1}$ and $V:=\int_{X}\theta^{n}>0$.
We recall that given $u_{0},u_{1}\in\mathcal{H}_{\theta}$, the weak geodesic from $u_{0}$ to $u_{1}$ is the path $t\mapsto u_{t}$ defined as follows : for any $x\in X$ and $z\in S:=\left\{0<\operatorname{Re}(z)<1\right\}$, setting $t=\operatorname{Re}(z)\in[0,1]$, we define
$u_{t}(x):=U(x,\operatorname{Re}(z)),$ (13)
where $U$ is defined as the supremum :
$U(x,z)=\sup\left\{V\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\theta})\,\Bigg|\,\limsup_{\operatorname{Re}(z)\rightarrow 0,1}V(x,z)\leq u_{0,1}(x)\right\}.$ (14)
Here $\pi_{1}:X\times S\rightarrow X$ is the projection on the first factor.
Proposition 3.1 ([18, Proposition 1.4]).
Let $U$ be defined as in (14). Then $U\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\theta})\cap L^{\infty}(X\times S)$. In addition $U$ satisfies, for any $(x,z,z^{\prime})\in X\times S\times S$ :
$\left|U(x,z)-U(x,z^{\prime})\right|\leq\left(\sup_{X}(u_{0}-u_{1})\right)\left|\operatorname{Re}(z)-\operatorname{Re}(z^{\prime})\right|.$
Finally, $U$ is continuous in $\operatorname{Amp}(\theta)\times\overline{S}$ and satisfies :
$\begin{cases}U|_{\left\{\operatorname{Re}(z)=0\right\}}=u_{0}\;;\quad U|_{\left\{\operatorname{Re}(z)=1\right\}}=u_{1}\\ \left(\pi_{1}^{*}\theta+dd^{c}_{(x,z)}U\right)^{n+1}=0,\end{cases}$ (15)
and it is the unique bounded $\theta$-plurisubharmonic solution to the system of equations (15).
In particular, for any $x\in X$, the map $t\mapsto u_{t}(x)$ defined by (13) is Lipschitz continuous. The Lipschitz constant being $\sup_{X}(u_{0}-u_{1})$, it is uniform in $x\in X$.
Moreover, the geodesic $U\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\theta})\cap L^{\infty}(X\times S)$ is the decreasing limit of geodesics (joining $u_{0},u_{1}$) with respect to $\omega_{\varepsilon}$. More precisely, if we consider
$U^{\varepsilon}:=\sup\left\{V\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\omega_{\varepsilon}})\;|\;\limsup_{\operatorname{Re}(z)\rightarrow 0,1}V(x,z)\leq u_{0,1}(x)\right\},$
then we have :
Proposition 3.2.
The sequence $\{U^{\varepsilon}\}$ is decreasing. In addition :
$U^{\varepsilon}\underset{\varepsilon\rightarrow 0}{\longrightarrow}U\quad\text{uniformly on }K\times\overline{S},$ (16)
for any compact subset $K\subset\operatorname{Amp}(\theta)$.
Démonstration.
The decreasing convergence on $X\times\overline{S}$ follows from [18, Proposition 1.6]. Let us prove that the convergence is uniform on any compact set $K\subset\operatorname{Amp}(\theta)$. Proposition 3.1 asserts that $U$ and $U^{\varepsilon}$ are continuous on $K\times\overline{S}$. Since the sequence $\{U^{\varepsilon}\}$ is decreasingly converging to $U$, Dini’s theorem gives uniform convergence on $K\times\overline{S}$. ∎
As before, for any $t\in[0,1]$ and $z\in\overline{S}$ such that $\operatorname{Re}(z)=t$ we set :
$u_{t}^{\varepsilon}(\cdot):=U^{\varepsilon}(\cdot,z).$ (17)
It follows from the previous proposition that for any compact subset $K\subset\operatorname{Amp}(\theta)$, we have the decreasing convergence of the weak $\omega_{\varepsilon}$-geodesics $t\mapsto u_{t}^{\varepsilon}$ towards the weak $\theta$-geodesic $t\mapsto u_{t}$ :
$\sup_{(x,t)\in K\times\overline{S}}\left|u^{\varepsilon}_{t}(x)-u_{t}(x)\right|\underset{\varepsilon\rightarrow 0}{\longrightarrow}0.$ (18)
We now assume that the forms $\theta$ and $\omega$ are both integral. Recall that a $2$-form is said to be integral if its cohomology class lies in the image $H^{2}(X,\mathbb{Z})_{b}$ of the natural map $H^{2}(M,\mathbb{Z})\rightarrow H^{2}_{dR}(M)$. We then give an approximation of the weak geodesics segments defined in (13) by paths coming from a finite dimensional space.
Let $t\mapsto u_{t}^{\ell}$ be the path defined in (17) with $\varepsilon=1/\ell$. Since $\omega_{\ell}$ is Kähler, we can consider smooth approximations as in [2, Theorem 1] : let $u_{0,j}$ (resp. $u_{1,j}$) be a decreasing sequence of smooth strictly $\omega_{\ell}$-plurisubharmonic functions decreasing towards $u_{0}$ (resp. $u_{1}$) and let $u^{\ell}_{t,j}$ be the weak geodesic between $u_{0,j}$ and $u_{1,j}$. Recall that the weak geodesics $u_{t}^{\ell}$ and $u_{t,j}^{\ell}$ are defined so that $u^{\ell}_{\operatorname{Re}(z)}(x)=U^{\ell}(x,z)$ and $u^{\ell}_{\operatorname{Re}(z),j}(x)=U^{\ell}_{j}(x,z)$ where
$U^{\ell}:=\sup\left\{V\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\omega_{\ell}})\;|\;\limsup_{\operatorname{Re}(z)\rightarrow 0,1}V(x,z)\leq u_{0,1}(x)\right\}$
$U^{\ell}_{j}:=\sup\left\{V\in\operatorname{PSH}(X\times S,\pi_{1}^{*}{\omega_{\ell}})\;\Bigg|\;\begin{cases}\limsup_{\operatorname{Re}(z)\rightarrow 0}V(x,z)\leq u_{0,j}(x)\\ \limsup_{\operatorname{Re}(z)\rightarrow 1}V(x,z)\leq u_{1,j}(x)\end{cases}\right\}.$
We then have the following :
Proposition 3.3.
Fix $\ell\in\mathbb{N}$. The sequence $\left(U^{\ell}_{j}\right)_{j\in\mathbb{N}}$ decreases to $U_{\ell}$ uniformly in $X\times\overline{S}$,
$\sup_{(x,t)\in X\times[0,1]}\left|u_{t,j}^{\ell}(x)-u_{t}^{\ell}(x)\right|\underset{j\rightarrow+\infty}{\longrightarrow}0.$ (19)
In particular we can choose $j=j(\ell)$ such that
$\sup_{(x,t)\in X\times[0,1]}\left|u_{t,j}^{\ell}(x)-u_{t}^{\ell}(x)\right|\leq\frac{1}{\ell}.$ (20)
The proof of such a result is somehow classical. We give details for reader’s convenience.
Démonstration.
It follows from the comparison principle that the sequence $\left(U_{j}^{\ell}\right)_{j\in\mathbb{N}}$ is decreasing. Let us show that this sequence converges towards $U^{\ell}$. First, since $U^{\ell}$ is a candidate in the supremum defining $U_{j}^{\ell}$ (Proposition 3.1), we have $U^{\ell}\leq U_{j}^{\ell},$ for any $j$. Passing to the limit for $j\rightarrow+\infty$ we get $U^{\ell}\leq W^{\ell}:=\lim_{j\rightarrow\infty}U_{j}^{\ell}$. In particular $W^{\ell}$ is a genuine $\pi_{1}^{*}\omega_{\ell}$-psh function on $M$. Hence to infer that $W^{\ell}\geq U^{\ell}$, we only need to check that $W^{\ell}$ has the right boundary conditions. Since $t\mapsto u_{t,j}^{\ell}$ is convex and $t\mapsto tu_{1,j}+(1-t)u_{0,j}$ is affine with same endpoints we have that for all $t\in[0,1]$,
$\quad u_{t,j}^{\ell}\leq tu_{1,j}+(1-t)u_{0,j}.$
Letting $j\rightarrow+\infty$ and taking $t=0,1$ give :
$\lim_{j\rightarrow\infty}u_{0,j}^{\ell}\leq u_{0}\;;\qquad\lim_{j\rightarrow\infty}u_{1,j}^{\ell}\leq u_{1}.$
In other words :
$\lim_{j\rightarrow\infty}U_{j}^{\ell}|_{\left\{\operatorname{Re}(z)=0\right\}}\leq u_{0}\;;\qquad\lim_{j\rightarrow\infty}U_{j}^{\ell}|_{\left\{\operatorname{Re}(z)=1\right\}}\leq u_{1}.$
This shows that $W^{\ell}$ is candidate for the $\sup$ defining $U^{\ell}$. Moreover, the convergence is uniform thanks to Dini’s Theorem : indeed, $\left(U_{j}^{\ell}\right)_{j\in\mathbb{N}}$ is a decreasing sequence and, since $\omega_{\ell}$ is Kähler, Proposition 3.1 asserts that all the functions $U_{j}^{\ell}$ and $U^{\ell}$ are continuous on $X\times\overline{S}$. ∎
Theorem 3.4.
Let $u_{0},u_{1}\in\mathcal{H}_{\theta}$ and $u_{t}$ be the weak geodesic joining them. There exists a sequence of path $(\hat{u}_{t,\ell})_{\ell\in\mathbb{N}}$ in the image of $\frac{1}{\ell}\operatorname{FS}_{k}$, where $k$ depends on $\ell$, such that for any compact $K\subset\operatorname{Amp}(\theta)$, we have :
$\sup_{(x,t)\in K\times[0,1]}\left|\hat{u}_{t,\ell}(x)-u_{t}(x)\right|\rightarrow_{\ell\rightarrow\infty}0.$
Démonstration.
We apply the quantization process to $t\mapsto u_{t,j}^{\ell}$, with $L$ being an ample line bundle whose curvature is $\ell\theta+\omega$. Hence
$\mathcal{H}_{\ell\theta+\omega}$$\hat{\mathcal{H}}(L^{k}).$$\scriptstyle{\operatorname{Hilb}_{k}}$$\scriptstyle{\operatorname{FS}_{k}}$ (21)
By construction $\ell u_{0,j}^{\ell},\ell u_{1,j}^{\ell}\in\mathcal{H}_{\ell\theta+\omega}$. We then consider the geodesic $0,1]\ni t\mapsto H_{t}^{k,j,\ell}\in\hat{\mathcal{H}}(L^{k})$ that joins $\operatorname{Hilb}_{k}(\ell u_{0,j}^{\ell})$ to $\operatorname{Hilb}_{k}(\ell u_{1,j}^{\ell})$. Theorem 2.2 ensures the existence of a constant $C_{\ell,j}$ such that
$\sup_{(x,t)\in X\times[0,1]}\left|\frac{1}{\ell}\operatorname{FS}_{k}\left(H^{k,j,\ell}_{t}\right)(x)-u_{t,j}^{\ell}(x)\right|\leq\frac{C_{\ell,j}}{\ell}\frac{\log k}{k}.$ (22)
We now choose $j=j(\ell)$ so that (20) holds, and since, for $\ell\in\mathbb{N}$ fixed, $\lim_{k\rightarrow\infty}\frac{C_{\ell,j(\ell)}}{\ell}\frac{\log k}{k}=0,$ we can choose $k=k(\ell)$ so that :
$\sup_{(x,t)\in X\times[0,1]}\left|\frac{1}{\ell}\operatorname{FS}_{k(\ell)}\left(H^{k(\ell),j(\ell),\ell}_{t}\right)(x)-u_{t,j(\ell)}^{\ell}(x)\right|\leq\frac{1}{\ell}$ (23)
[…]
$\sup_{(x,t)\in K\times[0,1]}\left|\hat{u}_{t,\ell}(x)-u_{t}(x)\right|$ $\leq\sup_{(x,t)\in K\times[0,1]}\left|\frac{1}{\ell}\operatorname{FS}_{k}\left(H^{k,j,\ell}_{t}\right)(x)-u_{t,j}^{\ell}(x)\right|$ (24)
$\qquad+\sup_{(x,t)\in K\times[0,1]}\left|u_{t,j}^{\ell}(x)-u_{t}^{\ell}(x)\right|$ (25)
$\qquad+\sup_{(x,t)\in K\times[0,1]}\left|u_{t}^{\ell}(x)-u_{t}(x)\right|$ (26)
$\leq\frac{1}{\ell}+\frac{1}{\ell}+\sup_{(x,t)\in K\times[0,1]}\left|u_{t}^{\ell}(x)-u_{t}(x)\right|.$ (27)
The conclusion follows from (18).
$u_{0}$$u_{1}$$u_{t}$$u_{t}^{\ell}$
Step 1
$u_{0}$$u_{0,j}$$u_{1}$$u_{1,j}$$u_{t}$$u_{t}^{\ell}$$u_{t,j}$
Step 2
$u_{0}$$u_{0,j}$$u_{1}$$u_{1,j}$$u_{0,j(\ell+1)}$$u_{1,j(\ell+1)}$$u_{0,j(\ell)}$$u_{1,j(\ell)}$$u_{t}$$u_{t,j}$$u_{t,j(\ell)}$$\hat{u}_{t,\ell}$$\hat{u}_{t,\ell+1}$
    Step 3
Figure 1: Quantization Process for Geodesics
4 Quantization in Sasaki geometry
In this section, we explain how to adapt the quantization of geodesics done in Section 3 to the Sasaki setting. Let $(M,\mathcal{S})$ be a Sasaki manifold. We assume $(M,\mathcal{S})$ to be regular or quasi-regular.
If $(M,\mathcal{S})$ is a regular Sasaki manifold over the Kähler manifold $(X,\omega)$, the space of potentials $\mathcal{H}(M,\mathcal{S})$ is homothetic to the space of Kähler potentials $\mathcal{H}_{\omega}$ (see [22]). Thus, the quantization of $\mathcal{H}(M,\mathcal{S})$ reduces to the quantization of $\mathcal{H}_{\omega}$.
We then consider the case $(M,\mathcal{S})$ to be a quasi-regular Sasaki manifold over the Kähler orbifold $(X,\omega)$. Let $\pi:M\rightarrow X$ be the quotient map and denote by $M^{\text{reg}}:=\pi^{-1}(X^{\text{reg}}),$ the set of points in $M$ whose orbit under the Reeb vector field is regular. By [boyer_sasakian_2007, Theorem 7.1.3 (iii)] we can assume $\omega$ to be an integral $2$-form.
We consider a resolution of singularities $\tilde{X}$ of the Kähler orbifold $(X,\omega)$ :
[…]
Here $\rho$ is a holomorphic map which induces a biholomorphism between $\rho^{-1}(X^{\text{reg}})$ and $X^{reg}$. By construction $\tilde{X}$ is a Kähler manifold endowed with a Kähler form $\tilde{\omega}$, while $\theta:=\rho^{*}\omega$ is a real closed $(1,1)$-form which is big and semi-positive. Observe that, since $X$ is projective, both $\tilde{\omega}$ and $\theta$ are integral forms. The above diagram induces :
$\mathcal{H}_{\theta}$$\mathcal{H}(M,\mathcal{S})$$\mathcal{H}_{\omega}.$$\scriptstyle{\mathrm{p}}$$\scriptstyle{\rho^{*}}$$\scriptstyle{\mathrm{f}}$ (29)
Here $\mathrm{f}$ is the correspondence between the set of basic functions on $M$ and the set of smooth functions on $X$. The map $\mathrm{f}$ is a 11-to-11 correspondence, and $\mathrm{p}$ is its inverse . Also, the set $\mathcal{H}_{\omega}$ is the set of smooth functions $u$ on the orbifold $X$ such that $\omega+dd^{c}u>0$ in the sense that the $2$-form $\omega+dd^{c}u$, computed on each local lifts of orbifold charts, is a positive $(1,1)$-form.
The quantization result in this setting states as follows :
Theorem 4.1.
Let $u_{0},u_{1}\in\mathcal{H}(M,\mathcal{S})$ and let $u_{t}$ be the weak geodesic joining $u_{0}$ and $u_{1}$. There exists a sequence of path $(\hat{u}_{t,\ell})_{\ell\in\mathbb{N}}$, which are image of geodesics in a finite dimensional space, such that for any compact set $K\subset M^{\text{reg}}$, we have the following uniform convergence :
$\sup_{(x,t)\in K\times[0,1]}\left|\hat{u}_{t,\ell}(x)-u_{t}(x)\right|\longrightarrow_{\lim_{\ell\rightarrow\infty}}0.$
Proof of Theorem 4.1.
We start observing that $v_{t}:=\rho^{*}\left(\mathrm{p}(u_{t})\right)$ is the weak geodesic in $\mathcal{H}_{\theta}$ joining $\rho^{*}\left(\mathrm{p}(u_{0})\right)$ and $\rho^{*}\left(\mathrm{p}(u_{1})\right)$. We then apply the quantization process (Theorem 3.4) to $v_{t}$. This gives a sequence of paths defined on $\tilde{X}$ :
$t\mapsto\hat{v}_{t,\ell}\;\in\mathcal{H}_{\theta+\frac{1}{\ell}\tilde{\omega}}$$\qquad\ell\in\mathbb{N}$.
Moreover, the sequence $\hat{v}_{t,\ell}$ converges uniformly (as $\ell\rightarrow+\infty$) to $v_{t}$ on each compact set $K\subset\operatorname{Amp}(\theta)=\rho^{-1}(X^{\text{reg}})$. Since $\rho$ is a biholomorphism on $\rho^{-1}(X^{\text{reg}})$, we infer that on any compact subset of $X^{\text{reg}}$, the following uniform convergence holds
$\rho_{*}\,\hat{v}_{t,\ell}\;\longrightarrow v_{t},\quad{\rm as}\;\,\ell\rightarrow+\infty.$
At the level of the Sasaki manifold $(M,\mathcal{S})$, the above ensures uniform convergence
$\mathrm{f}(\rho_{*}\,\hat{v}_{t,\ell})\longrightarrow u_{t}$
on any compact subset of $M^{\text{reg}}$. This ends the proof with $\hat{u}_{t,\ell}:=\mathrm{f}(\rho_{*}\,\hat{v}_{t,\ell}).$
Références
  • [1] B. Berndtsson. Probability measures associated to geodesics in the space of Kähler metrics. In Algebraic and Analytic Microlocal Analysis, pages 395–419. Springer, 2013.
  • [2] Z. Błocki and S. Kolodziej. On Regularization of Plurisubharmonic Functions on Manifolds. Proceedings of the American Mathematical Society, 135 :2089–2093, 2007.
  • [3] S. Boucksom. Divisorial Zariski decompositions on compact complex manifolds. In Annales scientifiques de l’Ecole normale supérieure, volume 37, pages 45–76, 2004.
  • [4] S. Boucksom. Variational and non-Archimedean aspects of the Yau–Tian–Donaldson conjecture. In Proceedings of the International Congress of Mathematicians : Rio de Janeiro 2018, pages 591–617. World Scientific, 2018.
  • [5] S. Boucksom, P. Eyssidieux, V. Guedj, and A. Zeriahi. Monge–Ampère equations in big cohomology classes. Acta mathematica, 205(2) :199–262, 2010.
  • [6] E. Calabi and X. Chen. The space of Kähler metrics II. Journal of Differential Geometry, 61(2) :173–193, 2002.
  • [7] X. Chen. The space of Kähler metrics. Journal of Differential Geometry, 56(2) :189–234, 2000.
  • [8] X. Chen. Space of Kähler metrics III – on the lower bound of the Calabi energy and geodesic distance. Inventiones mathematicae, 175 :453–503, 03 2009.
  • [9] X. Chen and S. Sun. Space of Kähler metrics (V)—Kähler quantization. In X. Dai and X. Rong, editors, Metric and Differential Geometry : The Jeff Cheeger Anniversary Volume, volume 297 of Progress in Mathematics, pages 19–41. Birkhäuser/Springer Basel, 2012.
  • [10] X. Chen and G. Tian. Geometry of Kähler metrics and foliations by holomorphic discs. Publ. Math. Inst. Hautes Etudes Sci., 2008.
  • [11] J. Chu, V. Tosatti, and B. Weinkove. $C^{1,1}$ regularity for degenerate complex Monge-Ampère equations and geodesic rays. Comm. Partial Differential Equations, 43(2) :292–312, 2018.
  • [12] T. Darvas. Morse theory and geodesics in the space of Kähler metrics. Proceedings of the American Mathematical Society, 142(8) :2775–2782, 2014.
  • [13] T. Darvas. The Mabuchi geometry of finite energy classes. Adv. Math., 285 :182–219, 2015.
  • [14] T. Darvas. Metric geometry of normal Kähler spaces, energy properness, and existence of canonical metrics. International Mathematics Research Notices, 2017(22) :6752–6777, 2017.
  • [15] T. Darvas. Geometric pluripotential theory on Kähler manifolds. In Advances in complex geometry, volume 735 of Contemp. Math., pages 1–104. Amer. Math. Soc., [Providence], RI, 2019.
  • [16] T. Darvas and L. Lempert. Weak geodesics in the space of Kähler metrics. Mathematical research letters, 19(5) :1127–1135, 2012.
  • [17] T. Darvas, H.-C. Lu, and Y. A. Rubinstein. Quantization in geometric pluripotential theory. Communications on Pure and Applied Mathematics, 73(5) :1100–1138, 2020.
  • [18] E. Di Nezza and V. Guedj. Geometry and topology of the space of Kähler metrics on singular varieties. Compositio Mathematica, 154(8) :1593–1632, 2018.
  • [19] S. K. Donaldson. Scalar curvature and projective embeddings, I. Journal of Differential Geometry, 59(3) :479–522, 2001.
  • [20] S. K. Donaldson. Scalar curvature and projective embeddings, II. The Quarterly Journal of Mathematics, 56(3) :345–356, 2005.
  • [21] S. Finski. Geometric quantization on big line bundles. 2025.
  • [22] T. Franzinetti. Isometries of the space of sasaki potentials. Annali della Scuola Normale Superiore di Pisa - Classe di Scienze, 13 :645–658, 2023.
  • [23] P. Guan and X. Zhang. A geodesic equation in the space of Sasakian metrics. Geometry and analysis, (1) :303–318, 2010.
  • [24] P. Guan and X. Zhang. Regularity of the geodesic equation in the space of Sasakian metrics. Advances in Mathematics, 230(1) :321–371, 2012.
  • [25] S. Helgason. Differential geometry, Lie groups, and symmetric spaces, volume 80 of Pure Appl. Math., Academic Press. Academic Press, New York, NY, 1978.
  • [26] L. Lempert. On the bergman kernels of holomorphic vector bundles. arXiv :2109.08593, 2021.
  • [27] L. Lempert and L. Vivas. Geodesics in the space of Kähler metrics. Duke Mathematical Journal, 162(7) :1369–1381, 2013.
  • [28] T. Mabuchi. Some symplectic geometry on compact Kähler manifolds. I. Osaka journal of mathematics, 24(2) :227–252, 1987.
  • [29] D. H. Phong and J. Sturm. The Monge-Ampère operator and geodesics in the space of Kähler potentials. Invent. Math., 166(1) :125–149, 2006.
  • [30] D. H. Phong and J. Sturm. Test configurations for k-stability and geodesic rays. Journal of Symplectic Geometry, 5(2) :221–247, 2007.
  • [31] J. Ross and D. Witt Nyström. Analytic test configurations and geodesic rays. Journal of Symplectic Geometry, 12(1) :125–169, 2014.
  • [32] J. Song and S. Zelditch. Bergman metrics and geodesics in the space of Kähler metrics on toric varieties. Analysis & PDE, 3(3) :295–358, 2010.